OpenStudy (anonymous):

pH of a substance is defined as pH = -log[H^+] where H^+ is the concentration of the Hydrogen Ion. The more concentrated it is, the more acidic the substance is. A pH of 7 is neutral, while a pH of 0 is extremely acidic. The pH of lemon juice is about 2.0, while a tomato juice is about 4.5 on the pH scale. How much more acidic (as in how much greater is the concentration of the Hydrogen Ion) is lemon juice than tomato juice?

OpenStudy (michele_laino):

if we apply the definition of \(pH\), we can write: \[\Large {\left[ {{{\text{H}}^ + }} \right]_{lemon}} = {10^{ - 2}}\] for lemon juice, and: \[\Large {\left[ {{{\text{H}}^ + }} \right]_{tomato}} = {10^{ - 4.5}} \simeq 3.16 \cdot {10^{ - 5}}\] for tomato juice

OpenStudy (michele_laino):

so we have this ratio: \[\Large \frac{{{{\left[ {{{\text{H}}^ + }} \right]}_{lemon}}}}{{{{\left[ {{{\text{H}}^ + }} \right]}_{tomato}}}} = \frac{{{{10}^{ - 2}}}}{{3.16 \cdot {{10}^{ - 5}}}} \simeq 3.16 \cdot {10^2}\]